OpenStudy (anonymous):

Expand The Logarithms : ln[ex^3/y^2z] I would really appreciate any explanation or help that could be given about this problem or like problem.

OpenStudy (anonymous):

\[\ln [ex ^{3}/y ^{2}z]\]

OpenStudy (anonymous):

ln(m/n)=ln(m)-ln(n)

OpenStudy (anonymous):

\[e ^{x}=1+x+x ^{2}/2!+x ^{3}/3!+.......\]

OpenStudy (anonymous):

I'm not really understanding in how you got there. Where are we getting the \[x ^{2}/2!\] from ? Also where exactly did the "ln" go ?

OpenStudy (anonymous):

You dont need that expansion. Just remember lne = 1. And ln( m . n) = lnm + ln(n) So the question becomes, \[lne + lnx^2 - lny^2 - lnz\]

OpenStudy (anonymous):

= 1 + 2lnx - 2lny - lnz.

OpenStudy (anonymous):

isnt it e^(x^3)?

OpenStudy (anonymous):

Not from the first comment.

OpenStudy (anonymous):

I'm understanding most all of it except the "+2lnx". from the original problem above ln[ex3/y2z] it shows an exponent of 3, does that get reduced or separated to become the "2lnx" ?

OpenStudy (anonymous):

it seems as if x^3 is in power and you are saying multiplication???????

OpenStudy (anonymous):

Oh that 3 should be an exponent not shown as being multiplied.

OpenStudy (anonymous):

lne+3inx-2lny-lnz

OpenStudy (anonymous):

So basically the exponent moves to the beginning of the sequence ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Alright, I understand now. Thank you very much for the help.

OpenStudy (anonymous):

\[\log _{b}(m ^{n})=n \log _{b}(m)\]